Solve
Solution for x: 3(x-6)+4=5(x+2)-6

Answer by stanbon(75887)

Solution for x:
3(x-6)+4=5(x+2)-6
---
3x-18+4 = 5x+10-6
----
3x-14 = 5x+4
---
2x = -18
---
x = -9
=============
Cheers,
Stan H.
=============